OpenStudy (anonymous):

Solve the exponential equation 125^(8x-2)=150 Help please?

OpenStudy (anonymous):

first take the log of both side

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

It doesn't say log tho. Isn't log the same as root 10?

OpenStudy (anonymous):

what do you mean it doesnt say log?

OpenStudy (anonymous):

The problem

OpenStudy (anonymous):

its an exponential equation

OpenStudy (anonymous):

log125^(8x-2)=log150 then (8x-2)log125=log150

OpenStudy (anonymous):

Can u combine the logs?

OpenStudy (anonymous):

not in this case

OpenStudy (anonymous):

to solve an exponential equation, you need to expand it not combine it.. and this case its already expanded.

OpenStudy (anonymous):

now distribute the left side (8x-2)log125=log150

OpenStudy (anonymous):

It would be by doing 10^125...right?

OpenStudy (anonymous):